OpenStudy (anonymous):

Factor: 9x^2 + 0x - 64 = 0

OpenStudy (anonymous):

does it really say \(0x\)??

OpenStudy (anonymous):

The original problem is just 9x^2 - 64 but my teacher told us to put +0x in between to make it easier.

OpenStudy (anonymous):

in any case \[9x^2-64\] is the difference of two square \[a^2-b^2=(a+b)(a-b)\] use \[a=3x,b=4\]

OpenStudy (anonymous):

really???? i would like to meet your teacher

OpenStudy (anonymous):

Lol okay, be my guest.

OpenStudy (anonymous):

is it clear what you need to do?

OpenStudy (anonymous):

no.. o.o Uh the way i get it is like example : (x - 9) (x - 4) = 0

OpenStudy (anonymous):

I need to find an answer like that

OpenStudy (anonymous):

yes but in this case you have "the difference of two squares" meaning you have \[a^2-b^2\] which factors as \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (lukecrayonz):

Satellite, honestly my teacher said to put that to, don't ask me why.. I never bothered

OpenStudy (lukecrayonz):

TI-89 After geometry, I got tired of factoring.. haha

OpenStudy (anonymous):

omg you don't get it, every problem i solve to get ex. (x-3) has x^2 in it..

OpenStudy (lukecrayonz):

Mira, do you know how to distribute..?

OpenStudy (anonymous):

i will write the answer \[9x^2-64=(3x+4)(3x-4)\] because \(9x^2\) is the square of \(3x\) and \(64\) is the square of \(4\)

OpenStudy (anonymous):

OH I GET IT THANKKSSS :D

OpenStudy (anonymous):

got that? we can do another like it \[16x^2-4=(4x+2)(4x-2)\]

OpenStudy (anonymous):

\[x^2-y^2=(x+y)(x-y)\] and so on

OpenStudy (lukecrayonz):

Satellite, fan me really quickly? Thanks